OpenStudy (anonymous):

for what values of f(x) = |x - 1| differentiable?

OpenStudy (freckles):

well think about f(x)=|x-1| for which value of x is there a sharp turn?

OpenStudy (freckles):

When you can name that value, then you have found the only value of x which f is not differentiable

OpenStudy (anonymous):

at 1

OpenStudy (anonymous):

what does that satement mean exactly?

OpenStudy (anonymous):

it is not differential at x=1

OpenStudy (anonymous):

\[\frac{ d }{ dx }\left| x-a \right|=\frac{ x-a }{ \left| x-a \right| }\frac{ d }{ dx }\left( x-a \right)\]